# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
Optical tweezer arrays have had a transformative impact on atomic and molecular physics over the past years, and they now form the backbone for a wide range of leading experiments in quantum computing, simulation, and metrology. Underlying this development is the simplicity of single particle control and detection inherent to the technique. Typical experiments trap tens to hundreds of atomic qubits, and very recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control. However, scaling to thousands of atomic qubits with long coherence times and low-loss, high-fidelity imaging is an outstanding challenge and critical for progress in quantum computing, simulation, and metrology, in particular, towards applications with quantum error correction. Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites while simultaneously surpassing state-of-the-art performance for several key metrics associated with fundamental limitations of the platform. Specifically, while scaling to such a large number of atoms, we also demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array. Further, we show trapping lifetimes close to 23 minutes in a room-temperature apparatus, enabling record-high imaging survival of $99.98952(1)\%$ in combination with an imaging fidelity of over $99.99\%$. Our results, together with other recent developments, indicate that universal quantum computing with ten thousand atomic qubits could be a near-term prospect. Furthermore, our work could pave the way for quantum simulation and metrology experiments with inherent single particle readout and positioning capabilities at a similar scale.
@@
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling control at the individual-particle level barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, quantum simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and optical frequency metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. They are part of broader efforts towards single atom control in optical traps, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2023, gyger_continuous_2024. For quantum computers, simulators of spin systems and tweezer clocks, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity. Such optically trapped atomic qubits coexist with other platforms that have single qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
@@
In particular, quantum error correction (QEC) requires large system sizes, while maintaining high fidelities and long coherence times, in order to scale up the number of logical qubits preskill_quantum_2018, google_quantum_ai_suppressing_2023. Even the most resource-efficient QEC protocols require at least several thousand physical qubits with an error-per-gate level of $10^{-3}$ to encode $>$100 error-corrected logical qubits bravyi_high-threshold_2023, xu_constant-overhead_2023. However, published results from universal quantum computing architectures based on ion traps and superconducting qubits currently leverage only tens to hundreds of qubits moses_race_2023, kim_evidence_2023, and with current technologies for fabrication and control, generally suffer from deleterious effects as they are scaled kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019.
@@
Neutral atoms in optical tweezer arrays are a promising platform towards rapid scalability in the near-term thanks to a programmable architecture that is readily adaptable to larger system sizes. In particular, universal quantum computing has recently been realized in such systems bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, and coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, alongside ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024 and erasure error detection scholl_erasure_2023, ma_high-fidelity_2023. Leveraging the versatility of the optical tweezer architecture, several applications have been shown, including executing quantum phase estimation graham_multi-qubit_2022, realizing the toric code structure bluvstein_quantum_2022, building an optical clock with ancilla-based readout finkelstein_universal_2024, and implementing logical qubit operations bluvstein_logical_2024.
In terms of current system sizes, tens to hundreds of qubits are controlled by typical tweezer array experiments young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2023, bluvstein_logical_2024, shaw_benchmarking_2024. Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024, gyger_continuous_2024, following earlier iterative reloading schemes singh_dual-element_2022, shaw_dark-state_2023; none of these experiments, however, report control of qubits or measurement of coherence times.
@@
Figure 1: Large-scale tweezer array. a, Representative single-shot image of single cesium atoms across a 12,001 tweezer array. Inset: magnified view of a subsection of the stochastically loaded array. b, Averaged image of single atoms across a 12,001 tweezer array. Inset: magnified view of a subsection of the averaged array. Atoms are spaced by ${\sim}7.15\,\mu$m and held in 1061 nm and 1055 nm optical tweezers. c, Schematic of the optical tweezer array generation. Tweezer arrays, generated by two spatial light modulators (SLM), at 1061 nm and 1055 nm are combined with orthogonal polarization, and focused through an objective with a numerical aperture (NA) of 0.65 and a field of view (FOV) of 1.5 mm in diameter. The $y$-axis in the figure corresponds to the direction of gravity. We collect scattered photons from single atoms through the same objective (NA$=0.55$ at the imaging wavelength) and image them on a qCMOS camera. d, Histogram of filling fraction. We load 6,139 single atoms on average per experimental iteration. This corresponds to loading 51.2% of the array on average, with a relative standard deviation of 1.13% over 16,000 iterations. e, Summary of the key metrics demonstrated in this work. The imaging survival quoted is a steady-state imaging survival probability, and the global single-qubit gate fidelity is obtained from global randomized benchmarking (RB), as detailed in the text.
@@
Here, we demonstrate a tweezer array with around 12,000 sites, that traps over 6,100 atomic qubits, simultaneously surpassing state-of-the-art values for metrics associated with fundamental limitations of the platform, including hyperfine qubit coherence time and trapping lifetime in a room-temperature apparatus, as well as combined imaging fidelity and survival (Fig. 1). Our results have implications for aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
@@
Our approach leverages high-power trapping of single atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext. Data Fig. 1a), enabling low-loss, high-fidelity imaging in combination with long hyperfine coherence times at the scale of 6,100 qubits (Fig. 1e). Specifically, we demonstrate imaging of single cesium-133 atoms with a survival probability of $99.98952(1)\%$, in combination with an imaging fidelity of $99.99374(8)\%$, surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019. This, alongside a 22.9(1) minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021 — over three times longer than previous records for tweezer arrays in room-temperature apparatuses covey_2000-times_2019 — provides realistic timescales for array operations in large scale arrays with minimal loss, e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
Importantly, we further demonstrate a coherence time of 12.6(1) s, a record for a hyperfine qubit tweezer array, surpassing previous values by almost an order of magnitude bluvstein_quantum_2022, graham_multi-qubit_2022 and approaching results for a single hyperfine qubit in a customized blue-detuned trap tian_coherence_2023, alkali atoms in an optical lattice dudin_light_2013, and nuclear qubits in a tweezer array barnes_assembly_2022. We also show a single-qubit gate fidelity of $99.9834(2)\%$ measured with global randomized benchmarking, limited by technical noise. We observe high uniformity across the sites in the array for loading probability (Ext. Data Fig. 2c), trap depth (Ext. Data Fig. 2d), imaging survival (Ext. Data Fig. 4a), scattering rate (Ext. Data Fig. 3d), and qubit frequency (Ext. Data Fig. 5d). Our results, when paired with other recent techniques for scaling single-site addressing zhang_scaled_2024, along with demonstrations of two-qubit gate operations evered_high-fidelity_2023, ma_high-fidelity_2023, finkelstein_universal_2024 and qubit rearrangement beugnon_two-dimensional_2007, bluvstein_quantum_2022, indicate that high-fidelity universal quantum computing with about ten thousand atomic qubits could be a near-term prospect, providing a path towards QEC with hundreds of logical qubits xu_constant-overhead_2023.
@@
We scale the optical tweezer array platform to around 12,000 tweezers, while prioritizing achieving long qubit coherence times and high fidelities (Fig. 1e). As such, we trap using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions to minimize hyperfine qubit decoherence from photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007. Cesium atoms possess the highest polarizability among the stable alkali metal atoms at near-infrared wavelengths where commercial fiber amplifiers routinely provide continuous-wave laser powers that exceed 100 W. Thus, a large number of traps can be generated with sufficient depth. A representative single shot image of the array is shown in Fig. 1a, and an averaged image is shown in Fig. 1b.
The atoms are spaced by ${\sim}7.15\,\mathrm{\mu m}$ and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth across the array (Methods). The tweezer light is combined with polarization and focused through a high numerical aperture objective with an unusually large field of view of 1.5 mm diameter that provides a large area for qubit trapping and manipulation in ultra-high vacuum (Fig. 1c). We measure an average trap depth of $k_{B}\times 0.20(1)$ mK, with a standard deviation of $11.4\%$ across all sites (Ext. Data Fig. 2d), enabling consistent loading probability per site.
@@
We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) beams that are parallel to the atom array plane ($x$-$z$ plane in Fig. 1c) for fluorescence imaging of single atoms, which also simultaneously cool the atoms. The photons scattered from each atom are collected on a quantitative CMOS (qCMOS) camera for 80 ms, using the same objective with which the traps are created. We find that each site is loaded with a single atom on average 51.2% of the time with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms in a large tweezer array (Ext. Data Fig. 2c). This allows us to load over 6,100 sites on average in each iteration (Fig. 1d).
@@
We detect and distinguish atomic presence in the array with high fidelity. Each image undergoes a binarization procedure (detailed in Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected). Note that we detect no more than one atom schlosser_sub-poissonian_2001 in each tweezer. We weight the collected photons in a $7\times 7$-pixel box centered around each site cooper_alkaline-earth_2018, so as to add more weight to pixels close to the center of each site’s point-spread function (Ext. Data Fig. 3a). The resulting signal is compared with a threshold to determine if an atom is present or not (Fig. 2).
@@
We characterize the imaging fidelity, defined as the probability of correctly labeling atomic presence in a site, with a model-free approach, where no assumption is made about the photon distribution from Fig. 2. To this end, we identify anomalous series of binary outputs norcia_microscopic_2018 in three consecutive images. For instance, $0\rightarrow 1\rightarrow 1$ would point to a false negative event in the first image, while $1\rightarrow 1\rightarrow 0$ could be due to atom loss during the second image or a false negative event in the third one. This approach allows us to precisely decouple inherent atom loss from false negatives or positives (Methods). From this we find an imaging fidelity of 99.99374(8)%. Crucial to this result are the homogeneous photon scattering rate across the array (Ext. Data Fig. 3d) and the consistency of the point-spread function even for sites at the edge of the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels). Furthermore, we find that treating each site separately with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%, indicating that the imaging parameters are sufficiently consistent across the atom array. Finally, we estimate that the imaging fidelity in the absence of atomic loss would be closer to 99.999% (Methods).
@@
The probability of not losing a single atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation. The vacuum-limited lifetime, in particular, quantifies a fundamental experimental limitation on the amount of time in which operations can be executed without loss of an atom. This can, for example, be applied to give an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
@@
We probe the vacuum-limited lifetime by using an empirically optimized cooling sequence consisting of a 10 ms 2D PGC cooling block every 2 s. By fitting the exponential decay of the atom survival, we find a $1/e$ lifetime of 22.9(1) min (Fig. 3a). This is more than three times longer than the longest reported lifetime for neutral atoms in an optical tweezer array using a room-temperature apparatus covey_2000-times_2019, and only within a factor of 5 of the longest reported lifetime in a cryogenic apparatus schymik_single_2021. The result indicates that the probability of losing a single atom across the entire array remains under 50% during 100 ms, a relevant timescale for dynamical array reconfiguration and quantum processor operation.
Figure 3: Long vacuum-limited lifetime and high imaging survival probability a, Vacuum-limited lifetime. Array-averaged survival fraction as a function of hold time is plotted. Two experiments are shown in the figure: with and without cooling. The green markers show data with a 10 ms 2D PGC block applied every 2 s (pulsed PGC), and the blue markers show the data without cooling (hold). The error bars are smaller than the markers. We find a $1/e$ lifetime of around 2.2 min without cooling. When the pulsed PGC block is applied, by fitting the data with $p(t)\propto\exp(-t/\tau)$, we find a vacuum lifetime of $\tau=22.9(1)$ min. b, Array-averaged survival fraction after many successive images. Between each image, we hold the atoms for 10 ms, without applying any cooling beams. We fit the data with $p(N)\propto p_{1}^{N}$, where $p(N)$ is the survival fraction after imaging $N$ times. From the fit, we find a steady-state imaging survival probability of $p_{1}=0.9998952(1)$. The light purple fill shows the estimated 67% confidence interval.

### v3

@@ Abstract
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have had a transformative impact on atomic and molecular physics over the past years, and they now form the backbone for a wide range of leading experiments in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. Underlying this development is the simplicity of single-particle control and detection inherent to the technique barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019. Typical experiments trap tens to hundreds of atomic qubits young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024, and very recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control pause_supercharged_2024, norcia_iterative_2024_mod, gyger_continuous_2024. However, scaling to thousands of atomic qubits with long coherence times and low-loss, high-fidelity imaging is an outstanding challenge and critical for progress in quantum computing, simulation, and metrology, in particular, towards applications with quantum error correction preskill_quantum_2018, xu_constant-overhead_2024. Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites while simultaneously surpassing state-of-the-art performance for several critical metrics that underpin the success of the platform. Specifically, while scaling to such a large number of atoms, we also demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array bluvstein_quantum_2022, graham_multi-qubit_2022. Further, we show trapping lifetimes close to 23 minutes in a room-temperature apparatus, enabling record-high imaging survival covey_2000-times_2019, blodgett_imaging_parsed_2023 of $99.98952(1)\%$ in combination with an imaging fidelity of over $99.99\%$. We lay out a detailed, near-term plan to enable zone-based quantum computing bluvstein_quantum_2022, bluvstein_logical_2024 with ${\sim}6,000$ atoms, and demonstrate a crucial ingredient thereof: coherent moves of up to $610\ \upmu$m with a fidelity of ${\sim}99.95\%$, as characterized through interleaved randomized benchmarking. Our results, together with other recent developments evered_high-fidelity_2023, ma_high-fidelity_2023, zhang_scaled_2024, finkelstein_universal_2024, indicate that universal quantum computing bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024 with ten thousand atomic qubits could be a near-term prospect. Furthermore, our work could pave the way for quantum simulation orourke_entanglement_2023, julia-farre_amorphous_2024 and metrology ludlow_optical_2015, rosenband_exponential_2013, haferkamp_closing_2020 experiments with inherent single particle readout and positioning capabilities at a similar scale.
@@
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling control at the individual-particle level barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, quantum simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and optical frequency metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. They are part of broader efforts towards single atom control in optical traps, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2024, gyger_continuous_2024. For quantum computers, simulators of spin systems and tweezer clocks, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity. Such optically trapped atomic qubits coexist with other platforms that have single-qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
@@
In particular, quantum error correction (QEC) requires large system sizes, while maintaining high fidelities and long coherence times, in order to scale up the number of logical qubits preskill_quantum_2018, google_quantum_ai_suppressing_2023. Even the most resource-efficient QEC protocols require at least several thousand physical qubits with an error-per-gate level of $10^{-3}$ to encode $>$100 error-corrected logical qubits bravyi_high-threshold_2024, xu_constant-overhead_2024. However, published results from universal quantum computing architectures based on ion traps and superconducting qubits currently leverage only tens to hundreds of qubits moses_race_2023, kim_evidence_2023, and with current technologies for fabrication and control, generally suffer from deleterious effects as they are scaled kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019.
@@
Neutral atoms in optical tweezer arrays are a promising platform towards rapid scalability in the near term thanks to a programmable architecture that is readily adaptable to larger system sizes. In particular, universal quantum computing has recently been realized in such systems bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, and coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, alongside ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024 and erasure error detection scholl_erasure_2023, ma_high-fidelity_2023. Leveraging the versatility of the optical tweezer architecture, several applications have been shown, including executing quantum phase estimation graham_multi-qubit_2022, realizing the toric code structure bluvstein_quantum_2022, building an optical clock with ancilla-based readout finkelstein_universal_2024, and implementing logical qubit operations bluvstein_logical_2024.
@@
Figure 1: Large-scale tweezer array. a, Representative single-shot image of single cesium atoms across a 11,998-site tweezer array. Inset: magnified view of a subsection of the stochastically loaded array. b, Averaged image of single atoms across a 11,998-site tweezer array. Inset: magnified view of a subsection of the averaged array. Atoms are spaced by $7.2\,\upmu$m and held in 1061 nm and 1055 nm optical tweezers. The contrast is enhanced for visual clarity. c, Schematic of the optical tweezer array generation. Tweezer arrays, generated by two spatial light modulators (SLM), at 1061 nm and 1055 nm are combined with orthogonal polarization, and focused through an objective with a numerical aperture (NA) of 0.65 and a field of view (FOV) of 1.5 mm in diameter. The $z$-axis in the figure is parallel to the direction of gravity. We collect scattered photons from single atoms through the same objective (NA$=0.55$ at the imaging wavelength) and image them on a qCMOS camera. d, Histogram of filling fraction. We load 6,139 single atoms on average per experimental iteration. This corresponds to loading 51.2% of the array on average, with a relative standard deviation of 1.13% over 16,000 iterations. e, Summary of the key metrics demonstrated in this work. The imaging survival quoted is a steady-state imaging survival probability, and the global single-qubit gate fidelity is obtained from global randomized benchmarking (RB) as detailed in the text.
In terms of current system sizes, tens to hundreds of qubits are controlled by typical tweezer array experiments young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024. Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024_mod, gyger_continuous_2024, following earlier iterative reloading schemes singh_dual-element_2022, shaw_dark-state_2023; none of these experiments, however, report control of qubits or measurement of coherence times.
@@
Here, we demonstrate a tweezer array with 11,998 sites, that traps over 6,100 atomic qubits, simultaneously surpassing state-of-the-art values for metrics underpinning the usefulness of the platform, including hyperfine qubit coherence time, trapping lifetime in a room-temperature apparatus, coherent transport distance and fidelity, as well as combined imaging fidelity and survival (Fig. 1). Our results have implications for aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
@@
Our approach leverages high-power trapping of single atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext. Data Fig. 1a), enabling low-loss, high-fidelity imaging in combination with long hyperfine coherence times at the scale of 6,100 qubits (Fig. 1e). Specifically, we demonstrate imaging of single cesium-133 atoms with a survival probability of $99.98952(1)\%$, in combination with an imaging fidelity of $99.99374(8)\%$, surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019. This, alongside a 22.9(1) minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021 — significantly longer than typical state-of-the-art vacuum lifetimes for tweezer arrays in room-temperature apparatuses— provides realistic timescales for array operations in large scale arrays with minimal loss, e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
Importantly, we further demonstrate a coherence time of 12.6(1) s, a record for a hyperfine qubit tweezer array, surpassing previous values by almost an order of magnitude bluvstein_quantum_2022, graham_multi-qubit_2022 and approaching results for a single hyperfine qubit in a customized blue-detuned trap tian_extending_2024, alkali atoms in an optical lattice wu_sterngerlach_2019, and nuclear qubits in a tweezer array barnes_assembly_2022. We also show a single-qubit gate fidelity of $99.9834(2)\%$ measured with global randomized benchmarking, limited by technical noise. We observe high uniformity across the sites in the array for loading probability (Ext. Data Fig. 2c), trap depth (Ext. Data Fig. 2d), imaging survival (Ext. Data Fig. 5a), scattering rate (Ext. Data Fig. 4d), and qubit frequency (Ext. Data Fig. 8a).
Finally, we lay out a realistic plan to implement zone-based universal quantum computation bluvstein_quantum_2022, bluvstein_logical_2024 with $\sim$6,000 atoms in the near term. We summarize in this work the key elements of our proposal, elucidated further in the Supplementary Information (SI). We notably demonstrate coherent atomic transport on a distance of 610 $\upmu$m with a fidelity of $\sim$99.95%, a crucial ingredient for scaling this architecture. Our results indicate that quantum computing with 6,000 atomic qubits is a near-term prospect, providing a path towards QEC with hundreds of logical qubits xu_constant-overhead_2024.
@@
We scale the optical tweezer array platform to 11,998 tweezers, while prioritizing achieving long qubit coherence times and high fidelities (Fig. 1e). As such, we trap using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions to minimize hyperfine qubit decoherence from photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007. Cesium atoms possess the highest polarizability among the stable alkali-metal atoms at near-infrared wavelengths where commercial fiber amplifiers routinely provide continuous-wave laser powers that exceed 100 W. Thus, a large number of traps can be generated with sufficient depth. A representative single shot image of the array is shown in Fig. 1a, and an averaged image is shown in Fig. 1b.
The atoms are spaced by ${\sim}7.2\,\mathrm{\upmu m}$ and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth across the array (Methods). The tweezer light is combined with polarization and focused through a high numerical aperture objective with an unusually large field of view of 1.5 mm diameter that provides a large area for atom trapping and manipulation in ultra-high vacuum (Fig. 1c). We measure an average trap depth of $k_{B}\times 0.18(2)$ mK, with a standard deviation of $11.4\%$ across all sites (Ext. Data Fig. 2d), enabling consistent loading probability per site.
Using Raman sideband spectroscopy with a 9.2 GHz amplitude-modulation technique levine_dispersive_2022, we measure radial and axial trapping frequencies of 29.3 kHz and 5.6 kHz, respectively (for the traps generated with the 1061 nm laser only; see Ext. Data Fig. 9). The setup used to address the atomic motion with a Raman configuration is detailed in the Methods. From this measurement we infer a $1/e^{2}$ tweezer waist $w_{0}=1.17(6)\ \mathrm{\upmu m}$.
@@
We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) beams that are approximately parallel to the atom array plane ($x$-$y$ plane in Fig. 1c) for fluorescence imaging of single atoms, which also simultaneously cool the atoms. The photons scattered from each atom are collected on a quantitative CMOS (qCMOS) camera for 80 ms, using the same objective with which the traps are created. We find that each site is loaded with a single atom on average 51.2% of the time with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms in a large tweezer array (Ext. Data Fig. 2c). This allows us to load over 6,100 sites on average in each iteration (Fig. 1d).
@@
We then detect and distinguish atomic presence in the array with high fidelity. Each image undergoes a binarization procedure (detailed in Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected). Note that we detect no more than one atom schlosser_sub-poissonian_2001 in each tweezer. We weight the collected photons in a $7\times 7$-pixel box centered around each site cooper_alkaline-earth_2018, so as to add more weight to pixels close to the center of each site’s point-spread function (Ext. Data Fig. 4a). The resulting signal is compared with a threshold to determine if an atom is present or not (Fig. 2).
@@
We characterize the imaging fidelity, defined as the probability of correctly labeling atomic presence in a site, with a model-free approach, where no assumption is made about the photon distribution from Fig. 2. To this end, we identify anomalous series of binary outputs norcia_microscopic_2018 in three consecutive images. For instance, $0\rightarrow 1\rightarrow 1$ would point to a false negative event in the first image, while $1\rightarrow 1\rightarrow 0$ could be due to atom loss during the second image or a false negative event in the third one. This approach allows us to precisely decouple inherent atom loss from false negatives or positives. From this we find an imaging fidelity of 99.99374(8)(note that we excise the first image, which we find has slightly lower fidelity and survival probability; Methods). Crucial to this result are the homogeneous photon scattering rate across the array (Ext. Data Fig. 4d) and the consistency of the point-spread function even for sites at the edge of the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels). Furthermore, we find that treating each site separately with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%, indicating that the imaging parameters are sufficiently consistent across the atom array. Finally, we estimate that the imaging fidelity in the absence of atomic loss would be closer to 99.999% (Methods).
Figure 3: Long vacuum-limited lifetime and high imaging survival probability a, Vacuum-limited lifetime. Array-averaged survival fraction as a function of hold time is plotted. Three experiments are shown in the figure: pulsed cooling, continuous cooling, and no cooling. The green markers show data with a 10 ms 2D PGC block applied every 2 s during the wait time (pulsed PGC). The red markers show data with 2D PGC block continuously applied during the wait time (continuous PGC), while the blue markers show the data without cooling during the wait time (no PGC). The error bars are smaller than the markers. We find a $1/e$ lifetime of around 2.2 min without cooling. When the pulsed PGC block is applied, by fitting the data with $p(t)\propto\exp(-t/\tau)$, we find a vacuum lifetime of $\tau=22.9(1)$ min. When the 2D PGC is applied continuously, we obtain $\tau=17.7(2)$ min. b, Array-averaged survival fraction after many successive images. Between each image, we hold the atoms for 10 ms, without applying any cooling beams. We fit the data with $p(N)\propto p_{1}^{N}$, where $p(N)$ is the survival fraction after imaging $N$ times. From the fit, we find a steady-state imaging survival probability of $p_{1}=0.9998952(1)$. The light purple fill shows the estimated 68% confidence interval.
@@
The probability of not losing a single atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation. The vacuum-limited lifetime, in particular, sets the ultimate upper bound on the duration during which operations can be executed without loss of an atom in a given experimental setup. This can, for example, be applied to give an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
@@
We probe the vacuum-limited lifetime by using an empirically optimized cooling sequence consisting of a 10 ms 2D PGC cooling block every 2 s. By fitting the exponential decay of the atom survival, we find a $1/e$ lifetime of 22.9(1) min (Fig. 3a). This is a significantly long timescale compared with state-of-the-art room-temperature atomic experiments, and only within a factor of five of the longest reported lifetime in a cryogenic apparatus schymik_single_2021. The result indicates that the probability of losing a single atom across the entire array remains under 50% during 100 ms, a relevant timescale for dynamical array reconfiguration and quantum processor operation.